CO2(g) + 2SO2(g) → CS(l) + 3O2(g)   ΔH = 1075 KJ


1
Expert's answer
2021-11-02T02:17:12-0400

"\Delta{H}_{reaction}=\sum{\Delta{H_{products}-\sum\Delta{H}_{reactants}}}=-393.5+2(-296.8)-87.9=-1075kJ/mol"


Answer: -1075 kJ/mol